# Large Language Model Tokenizer Bias: A Case Study and Solution on GPT-4o

### v2

@@ Abstract
Recent advancements in large language models (LLMs), such as GPT-4 and GPT-4o, have shown exceptional performance, especially in languages with abundant resources like English, thanks to extensive datasets that ensure robust training. Conversely, these models exhibit limitations when processing under-resourced languages such as Chinese and Korean, where issues including hallucinatory responses remain prevalent. This paper traces the roots of these disparities to the tokenization process inherent to these models. Specifically, it explores how the tokenizer’s vocabulary, often used to speed up the tokenization process and reduce tokens but constructed independently of the actual model training data, inadequately represents non-English languages. This misrepresentation results in the propagation of ’under-trained’ or ’untrained’ tokens, which perpetuate biases and pose serious concerns related to data security and ethical standards. We aim to dissect the tokenization mechanics of GPT-4o, illustrating how its simplified token-handling methods amplify these risks and offer strategic solutions to mitigate associated security and ethical issues. Through this study, we emphasize the critical need to rethink tokenization frameworks to foster more equitable and secure AI technologies.
@@ I Introduction
Byte Pair Encoding (BPE) has been widely adopted as a tokenization method due to its efficiency in handling extensive vocabularies and effectively managing rare words. LLMs, such as GPTs and Llama3, often employ a modified BPE tokenizer that ignores BPE merge rules for predefined vocabulary input tokens. GPT-4o shows a faster inference speed than GPT-4, attributing in part to increasing its built-in vocabulary size. However, this modification, aimed at optimizing performance, has inadvertently introduced a new challenge—biases in the tokenizer arising from the statistical properties of the training dataset result in critical hallucinations[1, 2]. Fig. 1 shows a case of how GPT-4 and GPT-4o handle a given Chinese text. Particularly, we observed that GPT-4o encounters specific issues when processing long Chinese tokens, while this phenomenon appears less in GPT-4, which employs a different tokenization approach. This discrepancy highlights the nuanced impact of tokenizers on model performance, particularly in the context of diverse linguistic inputs.
Figure 1: The exact tokenization process varies between GPT-4 and GPT-4o. Given the identical Chinese text, they use different tokenizers to process the input, leading to a variety of different tokens.
@@
In this paper, we delve into the inherent tokenization biases that affect LLMs, with a particular focus on GPT-4o. By conducting a comprehensive case study, we aim to elucidate the nuances of these biases and propose strategies to alleviate them. Our findings shed light on the critical need for unbiased tokenization approaches to enhance the robustness and fairness of future LLMs.
II Methodology
II-A Token Selection
@@ II-B Evaluation Criteria
II-B1 Sentence Generation Assessment
$N_{m}$ LLMs generate sentences using the designated token and its derivative shorter tokens (obtained by splitting the original token) for each token, with $2*N_{m}$ sentences generated.
@@ II-B2 Consistency
III Experiment
Figure 2: Sample sentences generated by GPT-4 and GPT-4o using long Chinese token from tiktoken-o200k_base (with the leading space removed) and its corresponding shorter tokens, along with human-assigned scores.
III-A Models
@@
III-B Experiment Setup
To assess the impact of token length in Chinese, according to Equation 1 with $N$ set to 20, a total of 166 tokens were sourced from the OpenAI’s tiktoken [^1] https://github.com/openai/tiktoken (o200k_base library, which GPT-4o especially used). As shown in Table I, the tokens range in size from 2 to 9, with each category sampled uniformly with 20 tokens except sizes 10 and 11. Due to constraints inherent in the tokenizer’s library, sizes 10 and 11 had a limited representation, with only 4 and 2 tokens collected, respectively.
@@ III-B Experiment Setup
III-C Experiment Results
III-C1 Tokens in Sentences
@@
III-C2 Sentences Ranking and Quality
Table III presents the rankings of sentences generated by GPT-4 and GPT-4o, evaluated by GPT-4 from privacy and security perspectives. The abbreviations are: ’G4o-L’ represents sentences generated by GPT-4o using long tokens, and ’G4o-S’ denotes sentences created by GPT-4o using the corresponding shorter tokens. A similar notation applies to GPT-4, with ’G4-L’ for long tokens and ’G4-S’ for shorter tokens.
Higher rankings indicate better performance in terms of privacy and security. Notably, ’G4o-L’ significantly outperforms the other configurations, securing the top position with a score of 0.6024, whereas ’G4-L’ ranks the lowest at 0.0783. Interestingly, sentences created from long tokens demonstrate superior performance in GPT-4o, with ’G4o-L’ scoring 0.6024 compared to ’G4o-S’ at 0.1205. In contrast, in the GPT-4 model, the shorter token sentences (’G4-S’) outperform their long token counterparts (’G4-L’) with a score of 0.1988 against 0.0783.
@@ III-C2 Sentences Ranking and Quality
Figure 3 displays the count of sentences rated as ”Score 5” across varying token sizes. The performance trends for G4o-S, G4-L, and G4-S are similar, tracking closely with the Total across all observed token sizes. This observation suggests that GPT-4 can effectively interpret longer token inputs and use them to generate coherent sentences. GPT-4 and GPT-4o exhibit comparable proficiency in dealing with their shorter counterparts, maintaining consistent quality in sentence construction.
However, there is a notable divergence in the performance of G4o-L, starting from a token size of 6. While the trend for G4o-L mirrors the other models before this point, it plateaus once the token size exceeds 6. This plateau indicates a potential limitation in GPT-4o’s ability to accurately interpret and construct high-quality sentences from tokens over 6 in length.
III-C3 Consistency
@@
IV Conclusion
This research has illuminated tokenization strategies’ profound effects on the performance and ethical integrity of LLMs, particularly GPT-4 and GPT-4o, and especially in their application to less-resourced languages such as Chinese and Korean. Our findings reveal that tokenization methods that prioritize operational efficiency—often at the cost of comprehensiveness—can severely undermine the models’ capabilities. These ’lazy’ tokenization approaches contribute not only to a degradation in linguistic accuracy and performance but also to heightened security risks and ethical challenges. Such issues arise because these methods do not adequately reflect the diversity of the training data, leading to biases that manifest as under-represented or entirely absent tokens.
In response, we have demonstrated that adopting more meticulous and comprehensive tokenization methods can significantly alleviate these issues. By ensuring that tokenizers capture a wider spectrum of linguistic nuances, we can mitigate the biases inherent in the training datasets. Moreover, our study advocates for an ongoing commitment to developing tokenization techniques that are both effective and ethical. Future research should explore innovative tokenization strategies that not only enhance model performance but also prioritize data security and ethical considerations. This involves rigorous testing across diverse linguistic datasets and continuous refinement of tokenizer algorithms to better handle the complexities of global languages.
Furthermore, the AI community must consider these tokenizer-related challenges as central to the responsible development of technology. By addressing these issues directly, we can advance toward creating LLMs that are not only technologically proficient but also uphold high ethical and security standards. This will ensure that as LLMs become increasingly integrated into global applications—from automated translation services to AI-driven content creation—they do so in a way that is inclusive, fair, and safe for all users.

### v3

# Problematic Tokens: Tokenizer Bias in Large Language Models

@@ Abstract
Recent advancements in large language models (LLMs), such as GPT-4 and GPT-4o, have shown exceptional performance, especially in languages with abundant resources like English, thanks to extensive datasets that ensure robust training. Conversely, these models exhibit limitations when processing under-resourced languages such as Chinese and Korean, where issues including hallucinatory responses remain prevalent. This paper traces the roots of these disparities to the tokenization process inherent to these models. Specifically, it explores how the tokenizer’s vocabulary, often used to speed up the tokenization process and reduce tokens but constructed independently of the actual model training data, inadequately represents non-English languages. This misrepresentation results in the propagation of ‘under-trained’ or ‘untrained’ tokens, which perpetuate biases and pose serious concerns related to data security and ethical standards. We aim to dissect the tokenization mechanics of GPT-4o, illustrating how its simplified token-handling methods amplify these risks and offer strategic solutions to mitigate associated security and ethical issues. Through this study, we emphasize the critical need to rethink tokenization frameworks to foster more equitable and secure AI technologies. The code and data are available at: https://github.com/yeyimilk/LLMGPT4o
@@ I Introduction
Byte Pair Encoding (BPE) has been widely adopted as a tokenization method due to its efficiency in handling extensive vocabularies and effectively managing rare words. LLMs, such as GPTs and Llama3, often employ a modified BPE tokenizer that ignores BPE merge rules for predefined vocabulary input tokens. GPT-4o shows a faster inference speed than GPT-4, attributing in part to increasing its built-in vocabulary size. However, this modification, aimed at optimizing performance, has inadvertently introduced a new challenge—biases in the tokenizer arising from the statistical properties of the training dataset result in critical hallucinations[1, 2]. Fig. 1 shows a case of how GPT-4 and GPT-4o handle a given Chinese text [^1] This variation is depicted in Figure 2, where ‘?’ indicates an encoding mismatch.. Particularly, we observed that GPT-4o encounters specific issues when processing long Chinese tokens, while this phenomenon appears less in GPT-4, which employs a different tokenization approach. This discrepancy highlights the nuanced impact of tokenizers on model performance, particularly in the context of diverse linguistic inputs.
Figure 1: The tokenization process differs between GPT-4 and GPT-4o, as they use distinct tokenizers for the same Chinese text, resulting in varying tokens. The English text was translated by human beings for reference.
@@
In this paper, we study the inherent tokenization biases that affect LLMs, with a particular focus on GPT-4o. By conducting a comprehensive case study, we aim to elucidate the nuances of these biases and propose strategies to alleviate them. Our findings shed light on the critical need for unbiased tokenization approaches to enhance the robustness and fairness of future LLMs.
II Motivating example
Figure 2: Motivating example. GPT-4o was not able to understand a common phrase used in Chinese scenario.
To illustrate the practical implications of tokenization biases, consider the scenario depicted in Figure 2. We asked GPT-4o the meaning of a Chinese phrase that translates to “WeChat official account: Win the lottery every day.” However, due to tokenization errors, GPT-4o might misinterpret this phrase as “WeChat money laundering” or “WeChat cash flow manipulation.” This misinterpretation stems from the model’s insufficient representation of such context-specific phrases in its training data. In reality, the phrase is a typical promotional message used in advertisements or notifications, not an indication of illicit activities.
Such a misinterpretation by automated systems, especially those that rely on tokenization for language understanding, highlights the challenges posed by ambiguous phrases and the dual meanings of words. If GPT-4o processes this phrase without understanding the correct context, it could tokenize and interpret the phrase solely in terms of financial manipulation. This misunderstanding could lead to the generation of inaccurate, misleading, or irrelevant content, potentially causing significant miscommunication in critical applications such as fraud detection, legal reviews, or financial reporting.
This example underscores the critical need for robust and context-aware tokenization methods. By improving how language models handle real-world, context-sensitive information, we can enhance their reliability and prevent errors that may have serious repercussions.
III Background or Related Work
Figure 3: The figure shows token samples from GPT-4o tokenizer, o200k_base, and their corresponding classification based on their content.
Figure 4: General workflow for determining problematic tokens.
Tokenization is a necessary step before feeding human input text to LLMs, which plays a crucial role in performance and efficiency.
III-A Tokenization
In the process of tokenization, each unique word, or small part, in the text sequence is identified and assigned a unique token ID. These tokenized outputs are mathematical entities and numerical values that represent the original input text in a new numeric format. Tokenization is fundamental to language model performance, with efficiency being directly dependent on the number of tokens that a given text is divided. As illustrated in Fig. 1, the same Chinese text is tokenized into 12 tokens by GPT-4, while GPT-4o tokenizes it into only 2 tokens. Generally, GPT-4o should exhibit better computational performance than GPT-4 when processing this text.
III-B BPE
Traditional tokenization techniques, such as using a word dictionary, are a significant challenge when confronted with out-of-vocabulary (OOV) issues. Hence, many LLMs include the strategy of employing sub-words, facilitating more comprehensive lexical support, including uncommon words, previously unseen words, and even different language lexica. The sub-word tokenization techniques include BPE, Word Piece Encoding, and Sentence Piece Encoding.
BPE is devised for data compression. In the context of NLP, BPE helps to mitigate the out-of-vocabulary (OOV) problem by representing an unknown word as a sequence of known sub-words. It starts with an initial vocabulary of individual characters. Then, it iteratively merges frequently adjacent symbols to form a new sub-word, thereby expanding the vocabulary.
The tokenizer in GPT models, including GPT-4 and GPT-4o, is a BPE model based on tiktoken. For GPT-4, the vocabulary size includes 100,256 predefined common tokens, while this number increases to 199,997 in GPT-4o. This tokenizer deviates from strict BPE merge rules when an input token is already part of the vocabulary. For instance, if no merge exists to produce ”hugging” and ”hugging” is part of the vocabulary, it will be returned as a single token rather than being split into smaller units like [”hug”, ”ging”]. Predefining tokens in this manner increases efficiency and accuracy by reducing the number of tokens that need to be processed, ultimately leading to faster and more precise language modeling.
III-C Tokenizer Bias in LLMs
While BPE and similar tokenization methods have their merits, they are not free from bias issues. This tokenizer bias arises from discrepancies in the statistical representation of vocabulary between the tokenizer and the LLM’s training corpus. Such disparities can result in numerous problems for LLMs[3]. Specifically, this misalignment may cause certain tokens to be underrepresented or completely absent during model training. When these ‘under-trained’ or ’untrained’ tokens—sometimes called ‘glitch tokens’[4, 5, 6]—are included in inputs[7], they can provoke unpredictable behaviors in the model, such as hallucinations or the production of nonsensical outputs[8]. The term ‘untrained’ is particularly used when there is definitive evidence that a token was not present in any training data for the model.
Fig. 3 illustrates Chinese token samples from the GPT-4o tokenizer, which can be categorized as related to gambling and adult content. Upon rigorous interpretation, these lexicographic entries seem peculiar, as they do not appear to be statistically significant common words or phrases worthy of being predefined in the tokenizer. This interpretation holds regardless of whether the comparison baseline is the general Chinese digital ecosystem or the training data source of GPT-3.
However, this phenomenon can be contextualized by considering the bifurcation of the Chinese internet environment into mainland China and regions beyond it. Most worthwhile and valuable internet data in mainland China is controlled by major Chinese corporations, which do not openly share their data and prohibit extraction through web crawlers. In contrast, regions beyond mainland China have a more open internet environment, hosting hundreds of thousands of gambling and adult websites. Given that the GPT-4o tokenizer may predominantly rely on open internet sources, these words could appear more frequently and be deemed statistically important within that dataset.
Nevertheless, these words should not be considered worthy of predefinition in the tokenizer because, in the broader context of the Chinese language—whether used on the internet or in daily communication—gambling and adult content terms are not commonly used. This indicates that biased or polluted data sources might influence the tokenizer’s vocabulary.
While tokenization significantly impacts the performance of LLMs [9], only a limited number of studies have explored this aspect in depth. When addressing bias, existing research predominantly focuses on areas such as training data, models, evaluation methodologies, and deployment strategies [10]. There is a noticeable gap in studies dedicated to identifying and analyzing biased tokens through model and tokenizer analysis, as well as examining problematic tokens using the model’s (un)embedding weights and tokenizer configuration[8, 5, 4, 9].
Our study addresses this gap by conducting a comprehensive analysis of extreme bias within tokenizers, particularly in the context of the Chinese language. We also propose methods to mitigate this bias and prevent hallucinations, thereby enhancing the reliability and fairness of LLMs in multilingual applications.
IV Methodology
Fig. 4 shows the workflow employed in our research. Candidate tokens are selected from the token library, followed by using different LLMs to understand those origin tokens and their corresponding generated smaller tokens. Those LLMs generated results were then evaluated by another LLM and human expert to determine the problematic tokens.
IV-A Token Selection
@@ IV-B Evaluation Criteria
IV-B1 Sentence Generation Assessment
$N_{m}$ LLMs generate sentences using the designated token and its derivative shorter tokens (obtained by splitting the original token) for each token, with $2*N_{m}$ sentences generated where $N_{m}$ denotes the number of LLMs.
@@ IV-B2 Consistency
V Experiment
Figure 5: Sample sentences generated by GPT-4 and GPT-4o using long Chinese token from tiktoken-o200k_base (with the leading space removed) and its corresponding shorter tokens, along with human-assigned scores. The English text was translated by human beings for reference.
V-A Models
@@
V-B Experiment Setup
All experiments described in this paper were completed as of August 11, 2024, ensuring consistent analysis under the model configurations available at that time.
To assess the impact5 of token length in Chinese, according to Equation 1 with $N$ set to 20, a total of 166 tokens were sourced from the OpenAI’s tiktoken [^2] https://github.com/openai/tiktoken (o200k_base library, which GPT-4o especially used). As shown in Table I, the tokens range in size from 2 to 9, with each category sampled uniformly with 20 tokens except sizes 10 and 11. Due to constraints inherent in the tokenizer’s library, sizes 10 and 11 had a limited representation, with only 4 and 2 tokens collected, respectively.
@@ V-B Experiment Setup
V-C Experiment Results
V-C1 Tokens in Sentences
@@
V-C2 Sentences Ranking and Quality
Table III presents the rankings of sentences generated by GPT-4 and GPT-4o, evaluated by GPT-4 from privacy and security perspectives. The abbreviations are: ‘G4o-L’ represents sentences generated by GPT-4o using long tokens, and ‘G4o-S’ denotes sentences created by GPT-4o using the corresponding shorter tokens. A similar notation applies to GPT-4, with ‘G4-L’ for long tokens and ‘G4-S’ for shorter tokens.
Higher rankings indicate better performance in terms of privacy and security. Notably, ‘G4o-L’ significantly outperforms the other configurations, securing the top position with a score of 0.6024, whereas ‘G4-L’ ranks the lowest at 0.0783. Interestingly, sentences created from long tokens demonstrate superior performance in GPT-4o, with ‘G4o-L’ scoring 0.6024 compared to ‘G4o-S’ at 0.1205. In contrast, in the GPT-4 model, the shorter token sentences (‘G4-S’) outperform their long token counterparts (‘G4-L’) with a score of 0.1988 against 0.0783.
@@ V-C2 Sentences Ranking and Quality
Figure 6 displays the count of sentences rated as “Score 5” across different token sizes. The performance trends for G4o-S, G4-L, and G4-S closely match the Total across all token sizes, indicating that GPT-4 can effectively handle longer token inputs to produce coherent sentences. Both GPT-4 and GPT-4o demonstrate similar skills in processing shorter tokens, consistently generating high-quality sentences.
However, G4o-L’s performance notably diverges starting at a token size of 6. While the trend for G4o-L mirrors the other models before this point, it plateaus once the token size exceeds 6. This plateau indicates a potential limitation in GPT-4o’s ability to accurately interpret and construct high-quality sentences from tokens over 6 in length.
V-C3 Consistency
@@
VI Discussion
The findings of this study provide valuable insights into the challenges and potential solutions for improving LLMs’ performance and ethical alignment when faced with data constraints. Our evaluation of GPT-4 and GPT-4o with long and short tokens highlights several key points:
VI-A Impact of Data Quality:
Relying on publicly available data, particularly in restrictive environments like China, introduces significant quality issues. The prevalence of low-quality content, such as gambling and pornography links in the training data, can lead to ethical misalignments and biases in the generated content. Our results show that GPT-4 generally outperforms GPT-4o in generating sentences that accurately incorporate both long and short tokens. However, the difference in performance between the two models narrows when shorter tokens are used, suggesting that token length plays a crucial role in model accuracy. Additionally, preliminary observations indicate that these abnormal tokens also negatively influence the output quality of GPT-3.5, further highlighting the importance of addressing data quality issues across different versions of LLMs.
VI-B Token Length and Model Performance:
The performance of GPT-4 and GPT-4o varies significantly with token length. GPT-4 demonstrates a superior capacity to construct sentences that accurately incorporate long tokens, while GPT-4o struggles more with long tokens, leading to lower performance and higher rates of unrelated sentence generation.
Recognizing that long tokens are often uncommon words and may be underrepresented in training data, we hypothesized that breaking long tokens into more common short tokens would enhance LLM understanding. Our results validate this hypothesis: both models exhibit improved performance when long tokens are segmented into shorter tokens, with GPT-4 maintaining a slight edge. This indicates that shorter tokens are easier for LLMs to handle accurately, which has implications for data preprocessing strategies.
VI-C Ethical and Security Considerations:
The evaluation of generated sentences from privacy and security perspectives revealed that sentences containing long tokens related to sensitive topics such as gaming and pornography scored lower. This highlights the importance of ethical data filtering to prevent such content from affecting the training and output of LLMs.
Interestingly, GPT-4o’s inability to accurately interpret long tokens may inadvertently result in sentences with fewer privacy and security concerns, as the generated content deviates from the intended sensitive contexts.
VI-D Mitigation Strategies:
Our study validates several mitigation strategies, including advanced data filtering techniques and ethical data collection practices. These strategies are crucial for enhancing LLMs’ robustness and ethical alignment, particularly in non-English contexts where data quality issues are more pronounced.
VII Conclusion
This study highlights the significant impact of tokenization strategies on the performance and ethical standards of LLMs, such as GPT-4 and GPT-4o, with a particular emphasis on their applications in less-resourced languages like Chinese. Our analysis has demonstrated that tokenization methods prioritizing operational efficiency over comprehensiveness can substantially compromise both linguistic accuracy and model integrity. These ’lazy’ approaches not only degrade performance but also exacerbate security risks and ethical issues by failing to represent the diversity of training data adequately, leading to significant biases.
A crucial practical outcome of our research is the creation of an open-source tool named ProblematicTokens[13], which aids in detecting and addressing tokenization flaws. This tool provides the AI community with essential capabilities to rigorously assess and enhance tokenization methods across varied linguistic contexts.
In closing, our findings emphasize the importance of meticulous tokenization in maintaining both the technical efficacy and ethical robustness of LLMs. We call for ongoing research and collaborative efforts to refine these practices, ensuring AI technologies operate equitably and securely. Addressing tokenizer-related challenges is vital for advancing the responsible development of LLMs, enabling their safe and fair integration into global applications, from automated translation to content creation.